# Coordinates from Context: Using LLMs to Ground Complex Location References

### v1

@@ Abstract
Geocoding is the task of linking a location reference to an actual geographic location and is essential for many downstream analyses of unstructured text. In this paper, we explore the challenging setting of geocoding compositional location references. Building on recent work demonstrating LLMs’ abilities to reason over geospatial data, we evaluate LLMs’ geospatial knowledge versus reasoning skills relevant to our task. Based on these insights, we propose an LLM-based strategy for geocoding compositional location references. We show that our approach improves performance for the task and that a relatively small fine-tuned LLM can achieve comparable performance with much larger off-the-shelf models.
@@ 1 Introduction
Extracting geospatial information from unstructured text is essential for many downstream analyses such as disaster response (Kumar and Singh, 2019), disease surveillance (Lee et al., 2013), and historical event analysis Tateosian et al. (2017). Geocoding is the task of linking a location reference to an actual geographic location, usually represented by a set of coordinates or an ID in a geographic database.
While there has been lots of work geocoding locations referred to with explicit names (e.g. "in Gaza City"), there has been little attention given to locations referred to with compositional descriptions involving other toponyms (e.g. "about midway between Winsford and Crewe") which occur frequently in domains that don’t have names for geographic regions of interest, e.g. environmental impact statements, food security outlooks Laparra and Bethard (2020). In such a setting, existing geocoding tools are not able to link the location being described by the entire phrase, only locations mentioned within the phrase Zhang and Bethard (2024). In addition, most geocoding tools can only link to locations already in a geographic database, which often do not include compositional locations.
@@
  • Based on these insights, we propose the first end-to-end strategy for geocoding compositional location references, leveraging both the geospatial reasoning capabilities of LLMs and the geospatial knowledge bases of traditional geoparsers. Furthermore, our novel use of bounding boxes [^1] A bounding box is a rectangular area defined by two longitudes and two latitudes, and written in the standard format of $\{\mathrm{lon}_{min},lat_{min},lon_{max},lat_{max}\}$ where $lon\in[-180,180]$ and $lat\in[-90,90]$. to link locations provides a simple and informative way to ground locations that don’t exist in a geographic database.
@@ 2 Related Work
Geocoding, also known as toponym resolution, aims to link some reference to a geographic entity to the correct reference in a target database or coordinate system Zhang and Bethard (2024). The majority of previous work has focused on geocoding location references that have an explicit location name, with difficult settings typically concerning ambiguous location names (e.g. "Paris" may refer to Paris, TX, USA or Paris, FR). Some researchers have looked at geocoding location references without explicit location names, such as indirect location references (e.g. "the Austrian capital") or nonstandard variants of location names (e.g. “Beantown” refers to Boston, MA, USA) Dredze et al. (2013); DeLozier et al. (2015); Kulkarni et al. (2021); Masis and O’Connor (2024).
@@
Very little attention has been given to geocoding complex or compositional location references, which do not contain an explicit location name and instead describe the location in relation to other toponyms. Thus far, there is only one dataset for this setting Laparra and Bethard (2020). The authors introduce five baseline methods for the task, although none using pretrained language models.
@@
Compositional location references must be linked to a coordinate system, not to a geographical database (e.g. GeoNames [^2] https://www.geonames.org or OpenStreetMap [^3] https://www.openstreetmap.org), since compositional locations often do not exist in geographic databases. While some prior work has examined geocoding without such databases DeLozier et al. (2015); Kulkarni et al. (2021); Sharma et al. (2023), they typically link locations to point coordinates which do not represent large or compositional geographic entities well. In our work, we link locations to a bounding box, which defines an area in terms of four sets of coordinates and allows us to ground locations much more precisely.
@@
Geoparsing is the broader task of both identifying and linking location references in unstructured text (Wang and Hu, 2019), essentially combining toponym recognition Hu et al. (2023a) and geocoding. As with geocoding, no geoparsing work to date has addressed compositional location references and methods primarily link location references to a geographic database or a single set of coordinates.
@@
LLMs’ geospatial knowledge and reasoning skills have begun to be evaluated in a growing body of work. (We define geospatial, following the ISO’s definition [^4] https://isotc211.geolexica.org/concepts/202/, as an ’implicit or explicit reference to a location relative to the Earth’ typically involving coordinates or addresses.) Regarding geospatial knowledge, recent work has suggested that LLMs learn and use an internal model of space which encodes locations’ approximate real-world geo-coordinates Gurnee and Tegmark (2024); Chen et al. (2023). Related studies have examined the disparities and differences in LLMs’ geospatial knowledge due to factors such as language used Faisal and Anastasopoulos (2023); Li et al. (2024) or model size Godey et al. (2024); Bhandari et al. (2023). To help explain how LLMs may acquire such knowledge, an examination of the Common Crawl Corpus – often used in the pretraining of LLMs – found that 19% of documents contained geospatial data and that a prominent source was Google Maps URLs, which frequently contain both coordinates and a natural language location name Ilyankou et al. (2024). However, to our knowledge no work has yet empirically evaluated LLMs as geospatial knowledge bases for downstream tasks.
@@
Regarding geospatial reasoning, recent work evaluating off-the-shelf LLMs has emphasized their promise but also their limitations for geospatial reasoning tasks Ji et al. (2025). Researchers have found that LLMs such as GPT-4o and Claude 3 Sonnet have limited success with reasoning over geospatial data, struggling to translate qualitative natural language terms to geospatial contexts O’Sullivan et al. (2024) and performing poorly at tasks like simple route planning Xu et al. (2024). Importantly, these studies show that LLMs have the capability to reason over geospatial information in both natural language and geo-coordinate formats.
@@
Using LLMs for geoparsing has recently been explored, with methods primarily using LLMs for the toponym recognition step Hu et al. (2024); Harrod et al. (2024); Xu et al. (2024); Hu et al. (2023b), to generate training data Yan and Lee (2024), or to embed natural language descriptions of point coordinate locations He et al. (2025). Thus far, LLMs have not been used for generating geospatial information or for reasoning-heavy settings.
@@ 3.1 Experimental Setup
Data. We use the GeoCoDe dataset, which contains location descriptions from English Wikipedia articles that are linked to the OpenStreetMap (OSM) geodatabase Laparra and Bethard (2020). For the current evaluation, we use GeoCoDe’s test set of 1,000 manually curated compositional location descriptions. Created by randomly sampling Wikipedia pages, the dataset includes locations from all continents.
Representing locations. By leveraging the OSM links in GeoCoDe, we retrieve gold geospatial information to represent each described location and location mentioned in a description. Locations can be represented as center coordinates, bounding boxes, or polygons, each with its own pros and cons. Center coordinates, which represent a location as a point, are used by the vast majority of geocoding datasets and are the simplest way to ground a location that does not exist in a geographic database. While adequate for point locations or small landmarks, they fall short when representing larger entities like states or countries. Polygons represent locations with an unlimited sequence of coordinates defining a closed shape, so are much more complex and thus rarely used; at present, there are no practical end-to-end geocoding systems using polygons. Bounding boxes represent a location as a rectangle-shaped area via two latitude-longitude coordinate pairs, efficiently encoding basic size and shape.
@@
Metrics. We use five metrics to evaluate predictions. Two are common geocoding metrics: coverage is the percentage of examples where the model produces a prediction, and distance error is the distance between the predicted and gold center coordinates or bounding box centroids. The other three metrics, introduced in Laparra and Bethard, measure the overlapping area between a predicted and gold bounding box: area precision is the fraction of overlapping area out of the entire predicted area, area recall is the fraction of overlapping area out of the entire gold area, and area F1-score is the harmonic mean of the precision and recall (calculated at test set-level, following prior work). Since these metrics measure area overlap, they are not used to evaluate center coordinate predictions.
@@ 3.2 Results
While previous work has argued that LLMs have some amount of geospatial knowledge, the evidence has mostly demonstrated learned connections between location names (usually cities) and their center coordinates Bhandari et al. (2023); Gurnee and Tegmark (2024). We explore LLMs’ geospatial knowledge with both center coordinates and bounding boxes, a more complex geo-entity representation, and our results caution against using LLMs as reliable geospatial knowledge bases in this setting. While the LLMs appear to perform slightly better when predicting a point than a bounding box, which may be due to point representations of geo-entities being a more common form of geospatial data in LLM pretraining data Ilyankou et al. (2024), both settings have relatively high average distance errors of over 200km.
On the other hand, our results show that LLMs perform promisingly on our geospatial reasoning baseline, with average distance errors of less than 35km. This supports prior work demonstrating that LLMs can reason over geospatial information encoded as both natural language and complex geometries Ji et al. (2025), even if complex geometries are not frequently in their pretraining data.
@@ 4.1 Task Definition
We formally define our task of geocoding compositional descriptions as follows. Each input to the geocoding system is a text description $d_{l}$ of a location $l$, where the set of locations mentioned in description $d_{l}$ is referred to as $M_{l}=\{m_{1},m_{2},...,m_{|M_{l}|}\}$. The description $d_{l}$ does not include the name of location $l$. The geocoding system is a combination of two functions. The first,
@@ 4.2 Our Approach
By augmenting with a traditional geoparser, we are able to add structured information from an external knowledge source to the LLM’s current context. This approach is related to the broader field of tool-augmented LLMs, in which LLMs leverage external tools Mialon et al. (2023). While traditional geoparsing tools can’t reason over compositional location descriptions or provide complex geospatial information (such as bounding boxes), they can link mentioned locations to a pair of coordinates. The mentioned locations’ coordinates can then serve as an accurate geospatially grounded starting point for the LLM to reason over where the described location is in relation to them.
In our experiments, we first simulate an oracle geoparser by retrieving locations’ center coordinates from the OSM database. This allows us to focus on the variation in performance introduced by LLMs, not by third-party geoparsing tools. We then evaluate a more realistic setup with a real geoparser as the recaller. For this, we use the Google Maps API [^6] https://developers.google.com/maps/documentation/geocoding which offers a popular geoparsing service broadly recognized as reliable Ozer et al. (2020).
@@ 4.3 Baselines
We also include results from the best performing method in prior work. GBSP: in this method, the reasoner is a grammar-based semantic parser and uses polygons of mentioned locations Laparra and Bethard (2020). We note that this method is not a practical end-to-end system since there is no recaller, and currently no system exists which can return polygons of mentioned locations.
@@ 5.1 Experimental setup
Models and training. We evaluate a range of state-of-the-art open-source models across sizes and from multiple families. For our methods that use an off-the-shelf model, we evaluate Llama3.1-8B-Instruct, Llama3.3-70B-Instruct, Llama3.1-405B-Instruct Dubey et al. (2024), Qwen3-14B-Instruct Yang et al. (2025), and Qwen2.5-72B-Instruct Yang et al. (2024). For methods that use a fine-tuned model, we evaluate Llama3.1-8B-Instruct and Qwen3-14B-Instruct. [^7] We evaluated Qwen3 models in the non-thinking mode.
@@ 5.2 Results
Comparing LLMs. Comparing model families, we see a similar pattern as in §3 where Qwen models (Table 3) perform slightly better when mentioned locations’ geospatial information is provided and Llama models (Table 2) have a slight edge when it is not. This suggests that Llama models may have better geospatial knowledge representations while Qwen models may have better geospatial reasoning abilities, which would support prior work demonstrating Qwen’s stronger reasoning abilities Wang et al. (2025). However, we note that disparities in performance are more dependent on model size than model family.
@@ 5.3 Qualitative analysis
First, a common error across model families was flipping the signs for latitude and/or longitude (e.g. "95.163" instead of "-95.163") in a predicted bounding box. We hypothesize that this error may occur due to how coordinates can represent distinctions between different hemispheres by either using positive/negative signs or cardinal directions. (e.g. "45°N" versus "45°S"). Both formats are widely used and both are likely present in LLM pretraining data corpora, so this may hinder LLMs from learning a consistent meaning for a negative or positive sign. We note that this error type occurred less with larger models and with the Geoparser-augmented approach, although fine-tuning did not seem to have much of an effect.
@@
This indicates that although providing LLMs with accurate geospatial information is generally helpful for our task, the LLMs may tend to overrely on the information without reasoning appropriately about its relationship or relevance to the desired output. This behavior is consistent with findings that LLMs perform much worse when irrelevant contextual information is included Shi et al. (2023) and that their ability to filter out irrelevant information decreases for more complex reasoning tasks Zhou et al. (2025). It may also be related to recent work demonstrating that reasoning models (e.g. DeepSeek-R1, Qwen3) have a strong tendency to defer to provided contextual knowledge, even when it contradicts parametric knowledge Marjanović et al. (2025).
@@ 6 Conclusion
In addition, our novel use of bounding boxes allows us to effectively ground locations that don’t exist in geographic databases and to evaluate methods with overlap-based metrics. Bounding boxes would likely be useful for other difficult geocoding settings, such as inherently vague geo-entities (e.g. "the Midwest") Jones et al. (2008). We encourage researchers to incorporate this way of representing locations into geocoding datasets and tools.
@@
While the domain we evaluate here is that of formal, impersonal writing (i.e. Wikipedia), we note that it is possible to apply geocoding methods to more personal writing (e.g. social media posts) and thus risk de-anonymizing online users through the inference of sensitive location information Kruspe et al. (2021); Dupre et al. (2022). For this reason, we have chosen to evaluate only open-source models in contrast to proprietary models, as they do not require sending potentially sensitive data to external servers and thus provide improved privacy. We also point to important recent work on geomasking techniques, which aim to protect the privacy of individuals while preserving spatial information in geodata Lorestani et al. (2024).

### v2

@@ Abstract
Geocoding is the task of linking a location reference to an actual geographic location and is essential for many downstream analyses of unstructured text. In this paper, we explore the challenging setting of geocoding compositional location references. Building on recent work demonstrating LLMs’ abilities to reason over geospatial data, we evaluate LLMs’ geospatial knowledge versus reasoning skills relevant to our task. Based on these insights, we propose an LLM-based strategy for geocoding compositional location references. We show that our approach improves performance for the task and that a relatively small fine-tuned LLM can achieve comparable performance with much larger off-the-shelf models. [^1] Our best-performing fine-tuned model is available at https://huggingface.co/tmasis/geocoding-complex-location-references
@@ 1 Introduction
Extracting geospatial information from unstructured text is essential for many downstream analyses such as disaster response (kumar2019location), disease surveillance (lee2013real), and historical event analysis tateosian2017tracking. Geocoding is the task of linking a location reference to an actual geographic location, usually represented by a set of coordinates or an ID in a geographic database.
While there is a large body of work on geocoding locations referred to with explicit names (e.g. "in Gaza City"), there has been little attention given to locations referred to with compositional descriptions involving other toponyms (e.g. "about midway between Winsford and Crewe") which occur frequently in domains that do not have names for geographic regions of interest, e.g. environmental impact statements, food security outlooks laparra2020dataset. In such a setting, existing geocoding tools are not able to link the location being described by the entire phrase, only locations mentioned within the phrase zhang2024survey. In addition, most geocoding tools can only link to locations already in a geographic database, which often do not include compositional locations.
@@
  • Based on these insights, we propose the first end-to-end strategy for geocoding compositional location references, leveraging both the geospatial reasoning capabilities of LLMs and the geospatial knowledge bases of traditional geoparsers. Furthermore, our novel use of bounding boxes [^2] A bounding box is a rectangular area defined by two longitudes and two latitudes, and written in the standard format of $\{\mathrm{lon}_{min},lat_{min},lon_{max},lat_{max}\}$ where $lon\in[-180,180]$ and $lat\in[-90,90]$. to link locations provides a simple and informative way to ground locations that do not exist in a geographic database.
@@ 2 Related Work
Geocoding, also known as toponym resolution, aims to link some reference to a geographic entity to the correct reference in a target database or coordinate system zhang2024survey. The majority of previous work has focused on geocoding location references that have an explicit location name, with difficult settings typically concerning ambiguous location names (e.g. "Paris" may refer to Paris, TX, USA or Paris, FR). Some researchers have looked at geocoding location references without explicit location names, such as indirect location references (e.g. "the Austrian capital") or nonstandard variants of location names (e.g. “Beantown” refers to Boston, MA, USA) dredze2013carmen; delozier2015gazetteer; kulkarni2021multi; masis2024earth.
@@
Very little attention has been given to geocoding complex or compositional location references, which do not contain an explicit location name and instead describe the location in relation to other toponyms al2019towards. Thus far, there is only one dataset for this setting laparra2020dataset. The authors introduce five baseline methods for the task, although none using pretrained language models.
@@
Compositional location references must be linked to a coordinate system, not to a geographical database (e.g. GeoNames [^3] https://www.geonames.org or OpenStreetMap [^4] https://www.openstreetmap.org), since compositional locations often do not exist in geographic databases. While some prior work has examined geocoding without such databases delozier2015gazetteer; kulkarni2021multi; sharma2024spatially, they typically link locations to point coordinates which do not represent large or compositional geographic entities well. In our work, we link locations to a bounding box, which defines an area in terms of four sets of coordinates and allows us to ground locations much more precisely.
@@
Geoparsing is the broader task of both identifying and linking location references in unstructured text (wang2019enhancing), essentially combining toponym recognition hu2023location and geocoding. As with geocoding, no geoparsing work to date has addressed compositional location references and methods primarily link location references to a geographic database or a single set of coordinates.
@@
LLMs’ geospatial knowledge and reasoning skills have begun to be evaluated in a growing body of work. (We define geospatial, following the ISO’s definition [^5] https://isotc211.geolexica.org/concepts/202/, as an ’implicit or explicit reference to a location relative to the Earth’ typically involving coordinates or addresses.) Regarding geospatial knowledge, recent work has suggested that LLMs learn and use an internal model of space which encodes locations’ approximate real-world geo-coordinates gurnee2024language; chen2023more. Related studies have examined the disparities and differences in LLMs’ geospatial knowledge due to factors such as language used faisal2023geographic; li2024land or model size godey2024scaling; bhandari2023large. To help explain how LLMs may acquire such knowledge, an examination of the Common Crawl Corpus – often used in the pretraining of LLMs – found that 19% of documents contained geospatial data and that a prominent source was Google Maps URLs, which frequently contain both coordinates and a natural language location name ilyankou2024quantifying. However, to our knowledge no work has yet empirically evaluated LLMs as geospatial knowledge bases for downstream tasks.
@@
Regarding geospatial reasoning, recent work evaluating off-the-shelf LLMs has emphasized their promise but also their limitations for geospatial reasoning tasks ji2025foundation. Researchers have found that LLMs such as GPT-4o and Claude 3 Sonnet have limited success with reasoning over geospatial data, struggling to translate qualitative natural language terms to geospatial contexts osullivan2024metric and performing poorly at tasks like simple route planning xu2024evaluating. Importantly, these studies show that LLMs have the capability to reason over geospatial information in both natural language and geo-coordinate formats.
@@
Using LLMs for geoparsing has recently been explored, with methods primarily using LLMs for the toponym recognition step hu2024toponym; harrod2024text; xu2024evaluating; hu2023geo, to generate training data yan2024georeasoner, or to embed natural language descriptions of point coordinate locations he2025geolocation. Thus far, LLMs have not been used for generating geospatial information or for reasoning-heavy settings.
@@ 3.1 Experimental Setup
Data. We use the GeoCoDe dataset [^6] https://github.com/EgoLaparra/geocode-data, which contains location descriptions from English Wikipedia articles that are linked to the OpenStreetMap (OSM) geodatabase laparra2020dataset. For the current evaluation, we use GeoCoDe’s test set of 1,000 manually curated compositional location descriptions. Created by randomly sampling Wikipedia pages, the dataset includes locations from all continents.
Representing locations. By leveraging the OSM links in GeoCoDe, we retrieve gold geospatial information to represent each described location and location mentioned in a description. Locations can be represented as center coordinates, bounding boxes, or polygons, each with its own pros and cons. Center coordinates, which represent a location as a point, are used by the vast majority of geocoding datasets and are the simplest way to ground a location that does not exist in a geographic database. While adequate for point locations or small landmarks, they fall short when representing larger entities such as states or countries. Polygons represent locations with an unlimited sequence of coordinates defining a closed shape, so are much more complex and thus rarely used; at present, there are no practical end-to-end geocoding systems using polygons. Bounding boxes represent a location as a rectangle-shaped area via two latitude-longitude coordinate pairs, efficiently encoding basic size and shape.
@@
Metrics. We use five metrics to evaluate predictions. Two are common geocoding metrics: coverage is the percentage of examples where the model produces a prediction, and distance error is the distance between the predicted and gold center coordinates or bounding box centroids. The other three metrics, introduced in laparra2020dataset, measure the overlapping area between a predicted and gold bounding box: area precision is the fraction of overlapping area out of the entire predicted area, area recall is the fraction of overlapping area out of the entire gold area, and area F1-score is the harmonic mean of the precision and recall (calculated at test set-level, following prior work). Since these metrics measure area overlap, they are not used to evaluate center coordinate predictions.
@@ 3.2 Results
While previous work has argued that LLMs have some amount of geospatial knowledge, the evidence has mostly demonstrated learned connections between location names (usually cities) and their center coordinates bhandari2023large; gurnee2024language. We explore LLMs’ geospatial knowledge with both center coordinates and bounding boxes, a more complex geo-entity representation, and our results caution against using LLMs as reliable geospatial knowledge bases in this setting. While the LLMs appear to perform slightly better when predicting a point than a bounding box, which may be due to point representations of geo-entities being a more common form of geospatial data in LLM pretraining data ilyankou2024quantifying, both settings have relatively high average distance errors of over 200km.
On the other hand, our results show that LLMs perform promisingly on our geospatial reasoning baseline, with average distance errors of less than 35km. This supports prior work demonstrating that LLMs can reason over geospatial information encoded as both natural language and complex geometries ji2025foundation, even if complex geometries are not frequently in their pretraining data.
@@ 4.1 Task Definition
We formally define our task of geocoding compositional descriptions as follows. Each input to the geocoding system is a text description $d_{l}$ of a location $l$, where the set of locations mentioned in description $d_{l}$ is referred to as $M_{l}=\{m_{1},m_{2},...,m_{|M_{l}|}\}$. The description $d_{l}$ does not include the name of location $l$.
The geocoding system is a combination of two functions (Figure 3). The first,
@@
Figure 3: We define our geocoding task as a combination of two functions. The first function (the recaller) retrieves mentioned locations’ geographical information, e.g. center coordinates. The second function (the reasoner) uses both the location description and mentioned location information to generate the described location’s bounding box.
@@ 4.2 Our Approach
By augmenting with a traditional geoparser, we are able to add structured information from an external knowledge source to the LLM’s current context. This approach is related to the broader field of tool-augmented LLMs, in which LLMs leverage external tools mialon2023augmented. While traditional geoparsing tools cannot reason over compositional location descriptions or provide complex geospatial information (such as bounding boxes), they can link mentioned locations to a pair of coordinates. The mentioned locations’ coordinates can then serve as an accurate geospatially grounded starting point for the LLM to reason over where the described location is in relation to them.
In our experiments, we first simulate an oracle geoparser by retrieving locations’ center coordinates from the OSM database. This allows us to focus on the variation in performance introduced by LLMs, not by third-party geoparsing tools. We then evaluate a more realistic setup with a real geoparser as the recaller. For this, we use the Google Maps API [^8] https://developers.google.com/maps/documentation/geocoding which offers a popular geoparsing service broadly recognized as reliable ozer2020creating.
@@ 4.3 Baselines
We also include results from the best performing method in prior work. GBSP: in this method, the reasoner is a grammar-based semantic parser and uses polygons of mentioned locations laparra2020dataset. We note that this method is not a practical end-to-end system since there is no recaller, and currently no system exists which can return polygons of mentioned locations.
@@ 5.1 Experimental setup
Models and training. We evaluate a range of state-of-the-art open-source models across sizes and from multiple families. For our methods that use an off-the-shelf model, we evaluate Llama3.1-8B-Instruct, Llama3.3-70B-Instruct, Llama3.1-405B-Instruct dubey2024llama, Qwen3-14B-Instruct yang2025qwen3, and Qwen2.5-72B-Instruct yang2024qwen2. For methods that use a fine-tuned model, we evaluate Llama3.1-8B-Instruct and Qwen3-14B-Instruct. [^9] We evaluated Qwen3 models in the non-thinking mode.
@@ 5.2 Results
Comparing LLMs. Comparing model families, we see a similar pattern as in §3 where Qwen models (Table 3) perform slightly better when mentioned locations’ geospatial information is provided and Llama models (Table 2) have a slight edge when it is not. This suggests that Llama models may have better geospatial knowledge representations while Qwen models may have better geospatial reasoning abilities, which would support prior work demonstrating Qwen’s stronger reasoning abilities wang2025octothinker. However, we note that disparities in performance are more dependent on model size than model family.
@@ 5.3 Qualitative analysis
First, an error across model families was flipping the signs for latitude and/or longitude (e.g. "95.163" instead of "-95.163") in a predicted bounding box. We hypothesize that this error may occur due to how coordinates can represent distinctions between different hemispheres by either using positive/negative signs or cardinal directions. (e.g. "45°N" versus "45°S"). Both formats are widely used and both are likely present in LLM pretraining data corpora, so this may hinder LLMs from learning a consistent meaning for a negative or positive sign. We note that this error type occurred less with larger models and with the Geoparser-augmented approach, and more with Qwen versus Llama models (see Table 10).
@@
This indicates that although providing LLMs with accurate geospatial information is generally helpful for our task, the LLMs may tend to overrely on the information without reasoning appropriately about its relationship or relevance to the desired output. This behavior is consistent with findings that LLMs perform much worse when irrelevant contextual information is included shi2023large and that their ability to filter out irrelevant information decreases for more complex reasoning tasks zhou2025gsm. It may also be related to recent work demonstrating that reasoning models (e.g. DeepSeek-R1, Qwen3) have a strong tendency to defer to provided contextual knowledge, even when it contradicts parametric knowledge marjanovic2025deepseek.
@@ 6 Conclusion
In addition, our novel use of bounding boxes allows us to effectively ground locations that do not exist in geographic databases and to evaluate methods with overlap-based metrics. Bounding boxes would likely be useful for other difficult geocoding settings, such as inherently vague geo-entities (e.g. "the Midwest") jones2008modelling. We encourage researchers to incorporate this way of representing locations into geocoding datasets and tools.
@@
While the domain we evaluate here is that of formal, impersonal writing (i.e. Wikipedia), we note that it is possible to apply geocoding methods to more personal writing (e.g. social media posts) and thus risk de-anonymizing online users through the inference of sensitive location information kruspe2021changes; dupre2022geospatial. For this reason, we have chosen to evaluate only open-source models in contrast to proprietary models, as they do not require sending potentially sensitive data to external servers and thus provide improved privacy. We also point to important recent work on geomasking techniques, which aim to protect the privacy of individuals while preserving spatial information in geodata lorestani2024privacy.